OpenStudy (anonymous):

Which of the following is the best starting point for approximating the square root of to the nearest whole number?

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

i would say the sq root of 4 because it is 2 and that is the lowest number of sq roots so i say sqrt 4

OpenStudy (anonymous):

what are your answer choices

OpenStudy (anonymous):

\[\sqrt{1}<\sqrt{73}<\sqrt{100} thats A \sqrt{64}<\sqrt{73}<\sqrt{81} thats B \sqrt{72}<\sqrt{73}\sqrt{74} thats C \sqrt{70}<\sqrt{73}<\sqrt{80} and thats D\]

OpenStudy (anonymous):

hold on a second i will work it out

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok i say the second set of numbers will i thinkd its B well the middle numbers

OpenStudy (anonymous):

ok ready for the next on

OpenStudy (anonymous):

yea